# Understanding Model Calibration - A gentle introduction and visual exploration of calibration and the expected calibration error (ECE)

### v2

@@ 2.1.1 Expected Calibration Error Drawbacks
The images of binning above provide a visual guide of how ECE could result in very different values if we used more bins or perhaps binned the same number of items instead of using equal bin widths. Such and more drawbacks of ECE have been highlighted by several works early on (Kumar et al., 2018; Nixon et al., 2019; Gupta et al., 2021; Zhang et al., 2020; Roelofs et al., 2022; Vaicenavicius et al., 2019; Widmann et al., 2019). However, despite the known weaknesses ECE is still widely used to evaluate confidence calibration in ML (Xiong et al., 2024; Yuan et al., 2024; Collins et al., 2023; Si et al., 2023; Mukhoti et al., 2023; Gao et al., 2024). This motivated this blogpost, with the idea to highlight the most frequently mentioned drawbacks of ECE visually and to provide a simple clarification on the development of different notions of calibration.
@@ 3.2 Binning Approach
One of the most frequently mentioned issues with ECE is its sensitivity to the change in binning (Kumar et al., 2018; Nixon et al., 2019; Gupta et al., 2021; Zhang et al., 2020; Roelofs et al., 2022). This is sometimes referred to as the Bias-Variance trade-off (Nixon et al., 2019; Zhang et al., 2020): Fewer bins reduce variance but increase bias, while more bins lead to sparsely populated bins increasing variance. If we look back to our ECE example with 9 samples and change the bins from 5 to 10 here too, we end up with the following:
@@
We can see that bin 8 and 9 each contain only a single sample and also that that half the bins now contain no samples. The above is only a toy example, however since modern models tend to have higher confidence values samples often end up in the last few bins (Naeini et al., 2015; Zhang et al., 2020), which means they get all the weight in ECE, while the average error for the empty bins contributes 0 to ECE.
@@ 3.3.2 Class-wise Calibration
To evaluate such different notions of calibration, some updates are made to ECE to calculate a class-wise error. One idea is to calculate the ECE for each class and then take the average (Nixon et al., 2019; Kull et al., 2019). Others, introduce the use of the KS-test for class-wise calibration (Gupta et al., 2021) and (Vaicenavicius et al., 2019) also suggest using statistical hypothesis tests instead of ECE based approaches.
@@ 3.3.3 Human Uncertainty Calibration
To evaluate human uncertainty calibration (Baan et al., 2022) introduce three new measures: the Human Entropy Calibration Error (EntCE), the Human Ranking Calibration Score (RankCS) and the Human Distribution Calibration Error (DistCE).

### v3

@@ 2.1.1 Expected Calibration Error Drawbacks
The images of binning above provide a visual guide of how ECE could result in very different values if we used more bins or perhaps binned the same number of items instead of using equal bin widths. Such and more drawbacks of ECE have been highlighted by several works early on (Kumar et al., 2018; Nixon et al., 2019; Gupta et al., 2021; Zhang et al., 2020; Roelofs et al., 2022; Vaicenavicius et al., 2019; Widmann et al., 2019; Futami & Fujisawa, 2024). However, despite the known weaknesses ECE is still widely used to evaluate confidence calibration in ML (Xiong et al., 2024; Yuan et al., 2024; Collins et al., 2023; Si et al., 2023; Mukhoti et al., 2023; Gao et al., 2024). This motivated this blogpost, with the idea to highlight the most frequently mentioned drawbacks of ECE visually and to provide a simple clarification on the development of different notions of calibration.
@@ 3.2 Binning Approach
One of the most frequently mentioned issues with ECE is its sensitivity to the change in binning (Kumar et al., 2018; Nixon et al., 2019; Gupta et al., 2021; Zhang et al., 2020; Roelofs et al., 2022; Famiglini et al., 2023). This is sometimes referred to as the Bias-Variance trade-off (Nixon et al., 2019; Zhang et al., 2020): Fewer bins reduce variance but increase bias, while more bins lead to sparsely populated bins increasing variance. If we look back to our ECE example with 9 samples and change the bins from 5 to 10 here too, we end up with the following:
@@
We can see that bin 8 and 9 each contain only a single sample and also that half the bins now contain no samples. The above is only a toy example, however since modern models tend to have higher confidence values samples often end up in the last few bins (Naeini et al., 2015; Zhang et al., 2020), which means they get all the weight in ECE, while the average error for the empty bins contributes 0 to ECE.
@@ 3.3.2 Class-wise Calibration
To evaluate such different notions of calibration, some updates are made to ECE to calculate a class-wise error. One idea is to calculate the ECE for each class and then take the average (Nixon et al., 2019; Kull et al., 2019). Another idea is to swap the L1-distance used in ECE with the L2 and use several ECE metrics to more effectively assess the overall level of calibration (Famiglini et al., 2023). Others, introduce the use of the KS-test for class-wise calibration (Gupta et al., 2021) and (Vaicenavicius et al., 2019) also suggest using statistical hypothesis tests instead of ECE based approaches. And other researchers develop a hypothesis test framework (TCal) to detect whether a model is significantly mis-calibrated (Lee et al., 2023) and Sun et al. (2024) build on this by developing confidence intervals for the L2-ECE.
@@ 3.3.3 Human Uncertainty Calibration
To evaluate human uncertainty calibration three new measures are introduced by Baan et al. (2022): the Human Entropy Calibration Error (EntCE), the Human Ranking Calibration Score (RankCS) and the Human Distribution Calibration Error (DistCE).